OpenStudy (anonymous):

solve for c f = 9/5c+32

OpenStudy (anonymous):

What do you mean by this? Do you have to write different equation?

OpenStudy (anonymous):

you basically have to get C by itself

OpenStudy (anonymous):

Oh, okay, I will try. Is it multiple choice?

OpenStudy (anonymous):

I KNOOOOW THIS ONEE

OpenStudy (anonymous):

its not multiply choice

OpenStudy (anonymous):

C=9/5(f-32)

OpenStudy (anonymous):

Ok...hmm...

OpenStudy (anonymous):

5/9(f-32)=C

OpenStudy (anonymous):

I'll give you a hint, This equation represents the relationship between Celsius temperatures and Fahrenheit temperatures. Currently it tells you what the Fahrenheit temp is. Just google for an equation that gives you the Celsius temp.

OpenStudy (anonymous):

c=5/9(f-32). you have to do the opposite. :)

OpenStudy (anonymous):

Does anyone know this one... 8.01], [8.02], [8.03], and [8.04] You are familiar with the following types of factoring: factoring out the Greatest Common Factor (GCF) factoring by grouping factoring trinomials of the form x2 + bx + c and ax2 + bx + c As you know, you need to know the first two types of factoring listed above in order to be successful in factoring trinomials of the form ax2 + bx + c. Part 1: In your own words, explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences. Part 2: If you were an Algebra 1 instructor and were creating a test on factoring trinomials of the form ax2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (anonymous):

whoa, which one is the correct answer

OpenStudy (anonymous):

P.S. Dwade03 is right! :)

OpenStudy (anonymous):

i got this by f=9/5c+32 *9 *9 9f=5c+32 /5 /5 9/5f=c+32 -32 +-32 9/5(f-32)=c or C=9/5(f-32) and i realized messed up sorry for that

OpenStudy (anonymous):

5/9(F-32)=C

OpenStudy (anonymous):

tah dahhh

OpenStudy (anonymous):

You have to switch it around, so it's the compete opposite of what it was earlier.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i am 2 you know and i proved it

OpenStudy (anonymous):

See...just do the opposite! And Dwade03, we hard you about 7 times already

OpenStudy (anonymous):

well theeen

OpenStudy (anonymous):

wtf this thing is messing up

OpenStudy (anonymous):

Your welcome! Good luck! :)